# Flex-Judge: Text-Only Reasoning Unleashes Zero-Shot Multimodal Evaluators

### v3

@@ 1 Introduction
Human-generated reward signals play a crucial role in both training and deploying generative models. They are commonly used to fine-tune models toward human-aligned behavior through preference optimization Rafailov et al. (2023); Meng et al. (2024) or reinforcement learning Ziegler et al. (2019); Stiennon et al. (2020). At inference time, they also guide inference-time decisions, e.g., best-of-$N$ selection Gui et al. (2024), output reranking Ma et al. (2023), or filtering based on quality or safety criteria, making them essential tools for test-time control. As models become more capable and are applied across diverse modalities and tasks, the need for high-quality, consistent human feedback continues to grow. However, scaling human feedback process is highly resource-intensive and challenging to generalize across different domains, highlighting a critical demand for more scalable and cost-effective alternatives that can reliably approximate human judgments Zheng et al. (2023); Wang et al. (2023c).
A promising alternative to manual feedback collection is to use large language models (LLMs) as proxy evaluators—an approach known as LLM-as-a-Judge Zheng et al. (2023). These models emulate human preferences via instruction-following prompts and have shown strong agreement with human ratings across tasks such as summarization Stiennon et al. (2020); Song et al. (2025) and dialogue Kim et al. (2024). In addition to reducing annotation costs, they can serve as reusable, modular evaluators and are often comparable to human judges in consistency. However, existing approaches are largely restricted to text-only scenarios Li et al. (2024b) and often require substantial amounts of paired preference data Kim et al. (2024) to generalize across evaluation types, such as single-score grading, pairwise comparison, or batch-level assessments.
Extending this paradigm to multimodal domains (e.g., vision-language generation or image caption ranking) presents unique challenges. Although some efforts have adapted LLM-as-a-Judge models into vision-language evaluators Chen et al. (2024a), they typically require extensive modality-specific annotations Xiong et al. (2024) and frequently fail to generalize across diverse formats without re-training or fine-tuning on each new task Lee et al. (2024). Moreover, the lack of publicly available multimodal preference datasets makes it difficult to systematically evaluate and train such models, often resulting in heavy reliance on proprietary models or benchmarks.
@@
  • Comparison with State-of-the-art (Section 3): We evaluate Flex-Judge on image, video, and audio reward benchmarks against commercial APIs Achiam et al. (2023); Team et al. (2023), large vanilla MLLMs, and open-source judges trained on costly multimodal datasets Lee et al. (2024); Xiong et al. (2024). Despite its simplicity and efficiency, Flex-Judge (7B model) outperforms open-source judges, even exceeding Gemini and GPT-4o on several MJ-Bench and GenAI-Bench subtasks. In-depth analyses are presented in Section 5.
  • Broader Impact (Section 4): We demonstrate real-world applications of Flex-Judge in the molecular domain by introducing Flex-Mol-LLaMA, the first judge model designed for molecular modality Kim et al. (2025). We showcase its utility in two key scenarios: (1) serving as a best-of-N selector for inference-time scaling, and (2) constructing training data for direct preference optimization (DPO) Rafailov et al. (2023). In both cases, reward-guided molecular MLLM achieves significant improvements, highlighting the practical solution in domains where modality-specific reward models are infeasible.
@@ 2.1 Motivation
Evaluating outputs across multiple modalities using foundation models is increasingly important, especially as generative models expand beyond language to include image, video, or audio Bai et al. (2023); Chu et al. (2024); Xu et al. (2025). While both proprietary LMs and open-source evaluators are widely used to assess the generative models’ response quality, two main challenges remain:
  • Concern of Commercial API: Issues regarding transparency, controllability, and affordability persist when utilizing proprietary LMs for evaluation tasks Kim et al. (2024). API model changes can silently degrade evaluation quality, raising concerns about the reliability of LLM-as-a-Judge with closed-source models Cai et al. (2025). For instance, Xiong et al. (2024) re-evaluated GPT-4V on the MLLM-as-a-Judge benchmark Chen et al. (2024a) and found a significant drop in evaluation performance.
  • Limited Support for Diverse Modalities: While judge models for language Kim et al. (2024) and vision-language tasks Lee et al. (2024); Xiong et al. (2024) have advanced significantly thanks to the availability of assessment data, evaluating other modalities, e.g., audio Chu et al. (2023, 2024), thermal heatmaps Yu et al. (2025), 3D point clouds Hong et al. (2023), and molecular structures Liu et al. (2024b), remains underexplored, with few effective judge models or publicly available training resources. For instance, evaluating a state-of-the-art molecular LLM Kim et al. (2025) often relies on GPT-4o to assess the soundness and relevance of its responses, which is not only hard to reproduce but also unreliable, as GPT-4o cannot handle molecular modalities.
@@
In multilingual LLMs Pires et al. (2019); Wu and Dredze (2019); Deshpande et al. (2022), it has been observed that fine-tuning on downstream tasks in one language can lead to performance improvements in other languages as well, demonstrating cross-lingual generalization. This suggests that when a shared representation space exists, task knowledge can effectively transfer across different languages.
@@ 2.2 Flex-Judge: Reasoning-Guided MLLM as a Judge
We begin by generating a high-quality “seed” textual dataset leveraging JudgeLRM Chen et al. (2025b), a reasoning-focused judge LM explicitly trained to evaluate AI responses with structured explanations. JudgeLRM uses crafted prompt templates to assess single or paired AI-generated outputs, producing detailed rationales enclosed within specialized tags (<think></think>). These reasoning annotations are detailed and comprehensive, explicitly addressing criteria such as correctness, completeness, consistency, relevance, and coherence.
@@
A critical advantage of our framework is the minimal data requirement. Specifically, we rely on only a 1K-sized corpus of high-quality textual reasoning annotations on text-only evaluation samples, making our approach highly cost-efficient, compared to MLLM judges such as Prometheus-Vision Lee et al. (2024) (150K image-text evaluation pairs for training) and LLaVA-Critic Xiong et al. (2024) (113K pairs). This corpus is carefully curated to exhibit the following properties:
@@
  • Quality & Difficulty: Following Muennighoff et al. (2025), we prioritize high-quality and high-difficulty samples when curating a training dataset for better sample efficiency. Specifically, we utilize JudgeLRM-7B to generate evaluation responses for prompts in the JudgeLM-100K dataset Zhu et al. (2025), and filter out responses whose predicted ratings mismatch with those annotated by GPT-4o in the original dataset. This quality-based selection process significantly improves the performance of the trained judge compared to random selection (i.e., blue; low-quality in 2(a)). Consistent with the findings of Muennighoff et al. (2025), we also observe that samples with longer reasoning chains (i.e., brown) improve judge performance across both language-only and multimodal settings. This observation supports the hypothesis that longer reasoning indicates higher difficulty, thereby enhancing the effectiveness of limited training data.
@@
  • Number of Data Samples & On-policy: We observe that training with a large number of samples can cause catastrophic forgetting Yang et al. (2024), where the LLM backbone exhibits diminished capacity to encode visual or audio features. As shown in 2(b), while JudgeLM Zhu et al. (2025) and PandaLM Wang et al. (2024b) performances improve with more training data, performances on multimodal benchmarks (MLLM-as-a-Judge and MJ-Bench) degrade, indicating a modality shift detrimental to multimodal understanding. We also find in 2(c) that lower-temperature decoding yields more effective training data, with lower initial losses. Since JudgeLRM-7B (the data generator) shares its LLM backbone with Flex-Judge, using these lower-loss, on-policy samples helps prevent catastrophic forgetting while preserving the language-side judge performance.
@@
  • Format Diversity: Compared to naïvely using unprocessed outputs from JudgeLRM-7B, which only supports pairwise scoring on a 1–10 scale as shown in Figure 7, we post-process the model’s outputs to support both single-score and pairwise grading, with scores mapped to either 1–10 or 1–5 scales depending on the instruction. Flex-Judge trained on these post-processed outputs demonstrates improved generalization to diverse evaluation formats, including a single-score grading and a batch-level ranking (i.e., where the number of response options exceeds three) as used in Chen et al. (2024a). Furthermore, we find that the post-processed variant is more robust to varied instruction styles, particularly when prompts emphasize different evaluation criteria across input pairs. The detailed results are provided in Section 3.2 (Table 1).
@@
Next, we use the reasoning seed dataset to fine-tune an MLLM, such as Qwen2.5-VL Bai et al. (2025) and Qwen2.5-Omni Xu et al. (2025). Despite being originally trained to handle both text and other modalities such as visual or audio inputs, the MLLM is fine-tuned exclusively on our textual reasoning annotations. We call this fine-tuned evaluator as Flex-Judge. Consequently, the structured and explicit reasoning provided by JudgeLRM enables Flex-Judge to learn how to systematically evaluate and justify preferences, significantly improving zero-shot transfer capabilities.
Figure 3: Reasoning process of Flex-Judge on the text-image alignment task (GenAI-Bench Li et al. (2024a)). Additional qualitative examples are found in Appendix D.
@@
At inference time, our model performs multimodal evaluation without additional fine-tuning or modality-specific annotations (i.e., training-free). We visualize an example output from Flex-Judge in Figure 3, demonstrating that the textual reasoning behind judgments has transferred across modalities. Furthermore, unlike most other evaluators, our Flex-Judge takes advantage of inference-time scaling to improve judgment performance by leveraging multiple reasoning paths, such as majority voting Wang et al. (2023b) or budget forcing Muennighoff et al. (2025), as addressed in Section 5.
@@ 3.1 Experimental Setup
Based on 1K-sized training dataset introduced in Section 2.2, we develop Flex-Omni-7B (image, video, and audio) and Flex-VL-7B (image and video) from Qwen2.5-Omni-7B Xu et al. (2025) and Qwen2.5-VL-7B Bai et al. (2025), respectively. We compare them against both commercial models with costly API usage Achiam et al. (2023); Team et al. (2023) and open-source models that require either extensive training data Lee et al. (2024); Xiong et al. (2024) or significantly more parameters Liu et al. (2024a); Wang et al. (2024a). For more implementation details, refer to Appendix B.
@@
To evaluate the quality of judge models, we measure how closely their assessments align with human annotations (or human-verified model evaluations Chen et al. (2024b); Li et al. (2024c)). The specific metric depends on the evaluation format: we measure (1) Pearson correlation Lee Rodgers and Nicewander (1988) for single-score grading tasks, (2) accuracy for pairwise (A/B) comparisons, and (3) normalized Levenshtein distance Levenshtein et al. (1966) for batch-level rankings (e.g., ABCD). Detailed descriptions for each benchmark and the evaluation prompts are provided in Appendix B.3.
@@ 3.2 Comparison with State-of-the-arts
Table 1: The overall performance of different MLLMs in judging, compared with human annotations on different datasets. We sample all judgments three times and average them to mitigate the bias. w. and w.o. tie represents tie and non-tie situations, respectively, by following Chen et al. (2024a). : reported results from LLaVA-Critic Xiong et al. (2024). : results from the original paper of MLLM-as-a-Judge Chen et al. (2024a). : Prometheus-Vision-13B Lee et al. (2024) is only trained under the Score setting, incapable of following Pair/Batch instructions. Training-free (TF) models have not been trained on multimodal evaluation data.
@@
Table 2: Comparison of MLLM evaluator performances on VL-RewardBench (Left) and MJ-Bench (Right). : results from the original works Li et al. (2024c); Chen et al. (2024b). Best and second best results.
@@
In Table 1, model judgments are compared with human ratings in MLLM-as-a-Judge benchmark. We compare Flex-Judge with the following model groups: commercial models including Gemini and GPT-4V (state-of-the-art closed-source models), and latest open-source baselines including Prometheus-Vision-7B Lee et al. (2024) and LLaVA-Critic-7B Xiong et al. (2024), trained on large-scale (>>100K) curated MLLM-as-a-judge datasets.
@@
Table 2 (right) presents the judge performance of generated images in MJ-Bench, whether they are well-aligned with a prompt, safe, or have artifacts. Flex-Judge models achieve higher scores than some commercial models (e.g., Claude 3 Opus) and all training-required judge models (PickScore-v1 Kirstain et al. (2023), HPS-v2.1 Wu et al. (2023), and ImageReward Xu et al. (2023)) by a large margin. While the baselines have high variance according to tasks, especially poor in safety check, our models perform highly consistent. The superiority of Flex-Judge in image generation assessment is further demonstrated in Table 3, where Flex-VL-7B with majority voting evaluation Wang et al. (2023b) outperforms GPT-4o and Gemini-1.5-Pro.
Table 3: Comparison of MLLM evaluator performance on GenAI-Bench. : results from the original work Li et al. (2024a).
@@
We further report judge performance on GenAI-Bench, including image generation, image editing, and video generation tasks, in Table 3. Notably, our Flex-VL-7B achieves the highest overall performance, as well as strong results in both image editing and video generation tasks, when inference-time scaling is applied Wang et al. (2023b). A detailed discussion of inference-time scaling is provided in Section 5. While Flex-Omni-7B shows lower performance initially, it also benefits significantly from inference-time scaling—unlike its non-reasoning baseline, Qwen2.5-VL-7B.
@@
Table 4 demonstrates the performance of our approach in speech quality evaluation, where the linear correlation coefficient (LCC) and Spearman’s rank correlation coefficient (SRCC) are calculated to assess the agreement between the model’s predicted scores and the human-annotated MOS and SS values. As observed in Chen et al. (2025a); Wang et al. (2025); Zezario et al. (2025), existing open-source audio LLMs like Qwen2-Audio struggle in quality assessment without specific fine-tuning, often exhibiting widespread hallucinations. While audio LLMs are primarily pretrained on semantics-related tasks like audio question answering (AQA) or captioning (AAC), they are less familiar with such qualitative evaluations, which are challenging due to different MOS standards depending on the dataset Wang et al. (2025). Still, Flex-Omni-7B outperforms all training-free judges and even Gemini-2.0-Flash.
Table 4: Audio MOS/SS prediction results on the test sets of the NISQA Mittag et al. (2021), BVCC Cooper and Yamagishi (2021), SOMOS Maniati et al. (2022), and VoxSim Ahn et al. (2024) datasets. System-level results are computed by averaging the utterance-level results within each text-to-speech system (not provided for NISQA due to the absence of system labels). : task-specific fine-tuning results from Wang et al. (2025).

### v4

@@ 1 Introduction
Human-generated reward signals play a crucial role in both training and deploying generative models. They are commonly used to fine-tune models toward human-aligned behavior through preference optimization [56, 48] or reinforcement learning [91, 61]. At inference time, they also guide inference-time decisions, e.g., best-of-$N$ selection [24], output reranking [43], or filtering based on quality or safety criteria, making them essential tools for test-time control. As models become more capable and are applied across diverse modalities and tasks, the need for high-quality, consistent human feedback continues to grow. However, scaling human feedback process is highly resource-intensive and challenging to generalize across different domains, highlighting a critical demand for more scalable and cost-effective alternatives that can reliably approximate human judgments [89, 71].
A promising alternative to manual feedback collection is to use large language models (LLMs) as proxy evaluators—an approach known as LLM-as-a-Judge [89]. These models emulate human preferences via instruction-following prompts and have shown strong agreement with human ratings across tasks such as summarization [61, 60] and dialogue [29]. In addition to reducing annotation costs, they can serve as reusable, modular evaluators and are often comparable to human judges in consistency. However, existing approaches are largely restricted to text-only scenarios [36] and often require substantial amounts of paired preference data [29] to generalize across evaluation types, such as single-score grading, pairwise comparison, or batch-level assessments.
Extending this paradigm to multimodal domains (e.g., vision-language generation or image caption ranking) presents unique challenges. Although some efforts have adapted LLM-as-a-Judge models into vision-language evaluators [11], they typically require extensive modality-specific annotations [76] and frequently fail to generalize across diverse formats without re-training or fine-tuning on each new task [32]. Moreover, the lack of publicly available multimodal preference datasets makes it difficult to systematically evaluate and train such models, often resulting in heavy reliance on proprietary models or benchmarks.
@@
  • Comparison with State-of-the-art (Section 3): We evaluate Flex-Judge on image, video, and audio reward benchmarks against commercial APIs [1, 62], large vanilla MLLMs, and open-source judges trained on costly multimodal datasets [32, 76]. Despite its simplicity and efficiency, Flex-Judge (7B model) outperforms open-source judges, even exceeding Gemini and GPT-4o on several MJ-Bench and GenAI-Bench subtasks. In-depth analyses are presented in Section 5.
  • Broader Impact (Section 4): We demonstrate real-world applications of Flex-Judge in the molecular domain by introducing Flex-Mol-LLaMA, the first judge model designed for molecular modality [27]. We showcase its utility in two key scenarios: (1) serving as a best-of-N selector for inference-time scaling, and (2) constructing training data for direct preference optimization (DPO) [56]. In both cases, reward-guided molecular MLLM achieves significant improvements, highlighting the practical solution in domains where modality-specific reward models are infeasible.
@@ 2.1 Motivation
Evaluating outputs across multiple modalities using foundation models is increasingly important, especially as generative models expand beyond language to include image, video, or audio [4, 16, 79]. While both proprietary LMs and open-source evaluators are widely used to assess the generative models’ response quality, two main challenges remain:
  • Concern of Commercial API: Issues regarding transparency, controllability, and affordability persist when utilizing proprietary LMs for evaluation tasks [29]. API model changes can silently degrade evaluation quality, raising concerns about the reliability of LLM-as-a-Judge with closed-source models [9]. For instance, Xiong et al. [76] re-evaluated GPT-4V on the MLLM-as-a-Judge benchmark [11] and found a significant drop in evaluation performance.
  • Limited Support for Diverse Modalities: While judge models for language [29] and vision-language tasks [32, 76] have advanced significantly thanks to the availability of assessment data, evaluating other modalities, e.g., audio [15, 16], thermal heatmaps [84], 3D point clouds [25], and molecular structures [41], remains underexplored, with few effective judge models or publicly available training resources. For instance, evaluating a state-of-the-art molecular LLM [27] often relies on GPT-4o to assess the soundness and relevance of its responses, which is not only hard to reproduce but also unreliable, as GPT-4o cannot handle molecular modalities.
@@
In multilingual LLMs [53, 74, 20], it has been observed that fine-tuning on downstream tasks in one language can lead to performance improvements in other languages as well, demonstrating cross-lingual generalization. This suggests that when a shared representation space exists, task knowledge can effectively transfer across different languages.
@@ 2.2 Flex-Judge: Reasoning-Guided MLLM as a Judge
We begin by generating a high-quality “seed” textual dataset leveraging JudgeLRM [12], a reasoning-focused judge LM explicitly trained to evaluate AI responses with structured explanations. JudgeLRM uses crafted prompt templates to assess single or paired AI-generated outputs, producing detailed rationales enclosed within specialized tags (<think></think>). These reasoning annotations are detailed and comprehensive, explicitly addressing criteria such as correctness, completeness, consistency, relevance, and coherence.
@@
A critical advantage of our framework is the minimal data requirement. Specifically, we rely on only a 1K-sized corpus of high-quality textual reasoning annotations on text-only evaluation samples, making our approach highly cost-efficient, compared to MLLM judges such as Prometheus-Vision [32] (150K image-text evaluation pairs for training) and LLaVA-Critic [76] (113K pairs). This corpus is carefully curated to exhibit the following properties:
@@
  • Quality & Difficulty: Following Muennighoff et al. [50], we prioritize high-quality and high-difficulty samples when curating a training dataset for better sample efficiency. Specifically, we utilize JudgeLRM-7B to generate evaluation responses for prompts in the JudgeLM-100K dataset [90], and filter out responses whose predicted ratings mismatch with those annotated by GPT-4o in the original dataset. This quality-based selection process significantly improves the performance of the trained judge compared to random selection (i.e., blue; low-quality in 2(a)). Consistent with the findings of [50], we also observe that samples with longer reasoning chains (i.e., brown) improve judge performance across both language-only and multimodal settings. This observation supports the hypothesis that longer reasoning indicates higher difficulty, thereby enhancing the effectiveness of limited training data.
@@
  • Number of Data Samples & On-policy: We observe that training with a large number of samples can cause catastrophic forgetting [81], where the LLM backbone exhibits diminished capacity to encode visual or audio features. As shown in 2(b), while JudgeLM [90] and PandaLM [70] performances improve with more training data, performances on multimodal benchmarks (MLLM-as-a-Judge and MJ-Bench) degrade, indicating a modality shift detrimental to multimodal understanding. We also find in 2(c) that lower-temperature decoding yields more effective training data, with lower initial losses. Since JudgeLRM-7B (the data generator) shares its LLM backbone with Flex-Judge, using these lower-loss, on-policy samples helps prevent catastrophic forgetting while preserving the language-side judge performance.
@@
  • Format Diversity: Compared to naïvely using unprocessed outputs from JudgeLRM-7B, which only supports pairwise scoring on a 1–10 scale as shown in Figure 7, we post-process the model’s outputs to support both single-score and pairwise grading, with scores mapped to either 1–10 or 1–5 scales depending on the instruction. Flex-Judge trained on these post-processed outputs demonstrates improved generalization to diverse evaluation formats, including a single-score grading and a batch-level ranking (i.e., where the number of response options exceeds three) as used in [11]. Furthermore, we find that the post-processed variant is more robust to varied instruction styles, particularly when prompts emphasize different evaluation criteria across input pairs. The detailed results are provided in Section 3.2 (Table 1).
@@
Next, we use the reasoning seed dataset to fine-tune an MLLM, such as Qwen2.5-VL [5] and Qwen2.5-Omni [79]. Despite being originally trained to handle both text and other modalities such as visual or audio inputs, the MLLM is fine-tuned exclusively on our textual reasoning annotations. We call this fine-tuned evaluator as Flex-Judge. Consequently, the structured and explicit reasoning provided by JudgeLRM enables Flex-Judge to learn how to systematically evaluate and justify preferences, significantly improving zero-shot transfer capabilities.
Figure 3: Reasoning process of Flex-Judge on the text-image alignment task (GenAI-Bench [35]). Additional qualitative examples are found in Appendix D.
@@
At inference time, our model performs multimodal evaluation without additional fine-tuning or modality-specific annotations (i.e., training-free). We visualize an example output from Flex-Judge in Figure 3, demonstrating that the textual reasoning behind judgments has transferred across modalities. Furthermore, unlike most other evaluators, our Flex-Judge takes advantage of inference-time scaling to improve judgment performance by leveraging multiple reasoning paths, such as majority voting [69] or budget forcing [50], as addressed in Section 5.
@@ 3.1 Experimental Setup
Based on 1K-sized training dataset introduced in Section 2.2, we develop Flex-Omni-7B (image, video, and audio) and Flex-VL-7B (image and video) from Qwen2.5-Omni-7B [79] and Qwen2.5-VL-7B [5], respectively. We compare them against both commercial models with costly API usage [1, 62] and open-source models that require either extensive training data [32, 76] or significantly more parameters [40, 67]. For more implementation details, refer to Appendix B.
@@
To evaluate the quality of judge models, we measure how closely their assessments align with human annotations (or human-verified model evaluations [14, 38]). The specific metric depends on the evaluation format: we measure (1) Pearson correlation [33] for single-score grading tasks, (2) accuracy for pairwise (A/B) comparisons, and (3) normalized Levenshtein distance [34] for batch-level rankings (e.g., ABCD). Detailed descriptions for each benchmark and the evaluation prompts are provided in Appendix B.3.
@@ 3.2 Comparison with State-of-the-arts
Table 1: The overall performance of different MLLMs in judging, compared with human annotations on different datasets. We sample all judgments three times and average them to mitigate the bias. w. and w.o. tie represents tie and non-tie situations, respectively, by following [11]. : reported results from LLaVA-Critic [76]. : results from the original paper of MLLM-as-a-Judge [11]. : Prometheus-Vision-13B [32] is only trained under the Score setting, incapable of following Pair/Batch instructions. Training-free (TF) models have not been trained on multimodal evaluation data.
@@
Table 2: Comparison of MLLM evaluator performances on VL-RewardBench (Left) and MJ-Bench (Right). : results from the original works [38, 14]. Best and second best results.
@@
In Table 1, model judgments are compared with human ratings in MLLM-as-a-Judge benchmark. We compare Flex-Judge with the following model groups: commercial models including Gemini and GPT-4V (state-of-the-art closed-source models), and latest open-source baselines including Prometheus-Vision-7B [32] and LLaVA-Critic-7B [76], trained on large-scale (>>100K) curated MLLM-as-a-judge datasets.
@@
Table 2 (right) presents the judge performance of generated images in MJ-Bench, whether they are well-aligned with a prompt, safe, or have artifacts. Flex-Judge models achieve higher scores than some commercial models (e.g., Claude 3 Opus) and all training-required judge models (PickScore-v1 [30], HPS-v2.1 [75], and ImageReward [78]) by a large margin. While the baselines have high variance according to tasks, especially poor in safety check, our models perform highly consistent. The superiority of Flex-Judge in image generation assessment is further demonstrated in Table 3, where Flex-VL-7B with majority voting evaluation [69] outperforms GPT-4o and Gemini-1.5-Pro.
Table 3: Comparison of MLLM evaluator performance on GenAI-Bench. : results from the original work [35].
@@
We further report judge performance on GenAI-Bench, including image generation, image editing, and video generation tasks, in Table 3. Notably, our Flex-VL-7B achieves the highest overall performance, as well as strong results in both image editing and video generation tasks, when inference-time scaling is applied [69]. A detailed discussion of inference-time scaling is provided in Section 5. While Flex-Omni-7B shows lower performance initially, it also benefits significantly from inference-time scaling—unlike its non-reasoning baseline, Qwen2.5-VL-7B.
@@
Table 4 demonstrates the performance of our approach in speech quality evaluation, where the linear correlation coefficient (LCC) and Spearman’s rank correlation coefficient (SRCC) are calculated to assess the agreement between the model’s predicted scores and the human-annotated MOS and SS values. As observed in [10, 68, 86], existing open-source audio LLMs like Qwen2-Audio struggle in quality assessment without specific fine-tuning, often exhibiting widespread hallucinations. While audio LLMs are primarily pretrained on semantics-related tasks like audio question answering (AQA) or captioning (AAC), they are less familiar with such qualitative evaluations, which are challenging due to different MOS standards depending on the dataset [68]. Still, Flex-Omni-7B outperforms all training-free judges and even Gemini-2.0-Flash.
Table 4: Audio MOS/SS prediction results on the test sets of the NISQA [49], BVCC [18], SOMOS [46], and VoxSim [2] datasets. System-level results are computed by averaging the utterance-level results within each text-to-speech system (not provided for NISQA due to the absence of system labels). : task-specific fine-tuning results from [68].